[figure: 9.7 cm]

THEORY OF THE SOLID-STATE PHYSICS ON THE TURN

R.J. Radwański^{a,b} and Z. Ropka^{a} aCenter for Solid State Physics, św. Filip 5, 31-150 Kraków bInst. of Physics, Pedagogical University, 30-084 Kraków.

The discovery of the high-temperature superconductivity in the copper oxide has revealed the enormous shortage of our general understanding of the 3d-ion compounds. Namely, the insulating state of La_{2}CuO_{4} contradicts the standard band-structure result that predicts it to be metal. This dramatic breakdown of the ordinary band-electron theory has been known already for years for 3d-ion monooxides. Different reparations do not lead to the consistent picture for 3d-ion systems (Mott insulators). At present after announcement of the essential importance of the single-ion effects and many-electron discrete states in description of compounds containing transition-metal atoms of the 3d and 4f groups the fundamental controversy has become the scientific fact. In the paper the modern solid-state physics theories will be discussed in connection with possibility for description of real systems like LaMnO_{3}, LaCoO_{3}, MgV_{2}O_{4}, NiO. It turns out that the remarkably consistent description has been obtained within the individualized-electron theory, that underlines the fact that 3d/4f ions constituting a solid preserve much of their individual quasi-atomic properties. We have found the existence of very strong correlations within the 3d shell that produce the highly-correlated 3d^{n} systems with large orbital magnetism. It is concluded that for the description of electronic and magnetic properties of real 3dion systems the fine electronic structure, resulting from crystal-field, spin-orbit coupling, local lattice distortions (Jahn-Teller effect) has to be taken into account.

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : R.J. Radwanski

Address for correspondence : Center for Solid State Physics, św. Filip 5, 31-150 Kraków

Email address : sfradwan@cyf-kr.edu.pl